# SuperRL: Reinforcement Learning with Supervision to Boost Language Model Reasoning

### v1

@@ 4.1 Experimental Setup
We evaluate the SuperRL framework across a diverse set of reasoning-focused benchmarks, covering both in-domain and cross-domain generalization scenarios.
First, we investigate the learning and generalization capabilities of each approach across different datasets and evaluation sets. The results reveal that pure RL excels in dense-reward environments, while Hybrid Actor achieves the best performance under sparse-reward environments, where both RL and sequential SFT+RL struggle. These findings delineate the effective operating regimes of each method and highlight the necessity of the Adaptive Switch.
Second, we demonstrate the generalizability of the Hybrid Actor design across multiple reinforcement learning algorithms (e.g., PPO and GRPO) and model scales (ranging from 1B to 8B). In most settings under sparse-reward conditions, hybrid yields better convergence, improved generalization, and more stable updates.
Third, we analyze the the Adaptive Switch module, showing how it accurately identifies reward density through early training probes and dynamically selects the appropriate training actor. Our results provide clear empirical support for its decision logic, ensuring optimal strategy selection without manual intervention.
Complete experimental settings, switch thresholds, and implementation details are provided in Appendix C.7 and Appendix C.8.
Table 1: Exact-match accuracy (%) across benchmarks using Qwen2.5-1.5B with GRPO. The RL baseline (gray) is compared to other methods, with deviations highlighted in green (gains) and red (drops); color intensity reflects magnitude. Rows are grouped by training dataset. Hybrid refers to the Hybrid Actor.
Training Method GSM8K MetaMath PRM12K LIMO OpenR1 HiTab
GSM8K RL 77.1 83.9 46.7 6.7 10.6 1.2
SFT 59.7 64.5 26.7 1.2 7.4 0.0
SFT+RL 74.7 77.3 35.1 6.7 9.8 0.0
Hybrid 77.0 79.9 14.0 2.6 14.1 1.7
MetaMath RL 74.0 81.7 50.8 5.0 12.5 6.0
SFT 70.1 76.5 33.9 2.4 7.9 0.0
SFT+RL 75.2 78.7 38.7 5.5 9.7 0.2
Hybrid 79.5 82.6 33.0 1.8 15.4 1.1
PRM12K RL 74.1 78.3 50.5 9.8 14.2 2.8
SFT 7.1 41.7 29.2 1.2 6.4 0.0
SFT+RL 69.2 74.3 47.7 8.5 13.2 0.0
Hybrid 72.0 74.8 50.3 11.0 15.7 3.3
OpenR1 RL 3.0 3.3 10.0 9.1 0.0 0.0
SFT 43.4 61.2 36.4 4.9 10.4 1.2
SFT+RL 69.7 72.9 46.5 9.1 15.7 2.1
Hybrid 70.0 72.7 46.0 9.8 15.0 10.4
LIMO RL 0.0 0.0 0.0 0.0 0.4 0.0
SFT 16.3 17.3 10.8 1.2 1.8 0.7
SFT+RL 28.7 48.7 19.3 3.0 3.2 0.0
Hybrid 70.0 72.7 47.2 6.9 7.6 0.1
HiTab RL 16.9 37.2 27.7 0.1 6.9 0.0
SFT 33.2 54.4 46.5 4.9 12.7 5.5
SFT+RL 35.7 36.6 30.1 9.1 14.0 60.7
Hybrid 44.1 41.2 34.7 5.5 8.8 64.2
4.2 RL Dominates in Dense-Reward Reasoning Tasks
Dense reward environments are characterized by frequent and informative feedback at each step of the generation process. This setting provides clear optimization signals, allowing models to efficiently explore the solution space without relying heavily on supervised demonstrations. Arithmetic and symbolic reasoning tasks with low output variability often fall into this category, making them well-suited for reinforcement learning. In Table 1, datasets such as GSM8K, MetaMathQA, and PRM12K exemplify dense reward settings. Across these benchmarks, RL demonstrates strong performance, highlighting its ability to exploit dense supervision for effective learning and generalization.
RL demonstrates strong performance under dense reward settings, achieving the highest accuracy on in-domain tasks and exhibiting superior generalization compared to all other methods. Concretely, when trained on GSM8K, RL achieves an exact-match accuracy of 77.1%, outperforming other methods. A similar pattern holds on MetaMathQA, where RL reaches 81.7%, surpassing SFT (76.5%) and SFT+RL (78.7%). These results highlight RL’s ability to effectively leverage dense and well-aligned reward signals. Moreover, RL exhibits strong cross-task generalization. When trained on GSM8K and evaluated on PRM12K, RL attains an accuracy of 46.7%, significantly outperforming SFT (26.7%), SFT+RL (35.1%), and the Hybrid method (14.0%). The pattern can be extended to all the three datasets metioned above. This suggests that RL-learned policies capture transferable reasoning patterns that extend beyond the source task.
4.3 Hybrid Actor Thrives in Sparse-Reward Reasoning Tasks
Table 2: KL-loss range and variance comparison: RL vs Hybrid Actor
Dataset KL Range (RL) KL Range (Hybrid) Max $\Delta$ KL Var (RL) KL Var (Hybrid) Var $\Delta$
MetaMath [2.63e-4, 0.6055] [8.59e-4, 0.2732] -54.9% 0.009388 0.002664 -71.6%
GSM8K [7.93e-4, 0.5306] [7.17e-4, 0.2770] -47.8% 0.007579 0.002475 -67.3%
PRM12K [5.68e-4, 0.2616] [7.96e-4, 0.0900] -66.0% 0.002159 0.000379 -82.4%
LIMO [3.60e-4, 0.7732] [3.59e-4, 0.2391] -69.1% 0.02489 0.001494 -94.0%
OpenR1 [3.01e-4, 1.9110] [6.78e-4, 0.0952] -95.0% 0.01637 0.000239 -98.5%
HiTab [7.58e-4, 1.9600] [6.59e-4, 1.0387] -47.0% 0.006257 0.01079 +72.4%
Sparse reward settings are characterized by infrequent or weak reward signals, often posing significant challenges to exploration and optimization. Datasets such as LIMO, OpenR1, and HiTab exemplify this regime.
In these environments, hybrid consistently outperforms other methods. For instance, on HiTab, where RL fails to learn (0.0%), the hybrid model achieves 64.2% accuracy—exceeding SFT+RL (60.7%) and SFT (5.5%). Similar gains are observed on OpenR1 and LIMO, where the hybrid method retains strong performance despite limited supervision. Notably, this robustness extends beyond symbolic reasoning tasks to semi-structured domains like TableQA dataset HiTab, highlighting the Hybrid Actor’s ability to generalize across diverse reasoning tasks.
Beyond accuracy, Hybrid Actor enhances both learning stability and generalization, especially when compared to the SFT+RL method. Unlike SFT+RL, which suffers from abrupt shifts between supervised and reinforcement phases, Hybrid Actor maintains a continuous inductive prior, leading to smoother updates and more consistent optimization. As shown in Table 2, Hybrid Actor significantly reduces KL-loss variance across challenging tasks, achieving reductions of 71.6% on MetaMath, 67.3% on GSM8K, and up to 98.5% on OpenR1. These reductions indicate fewer erratic updates and greater training stability. On HiTab, the KL variance under Hybrid Actor is higher (+72.4%)—not due to instability, but because reward signals in this dataset remain consistently zero,
These stability benefits also lead to stronger generalization across diverse reasoning tasks. For example, when trained on LIMO—a sparse-reward dataset where vanilla RL completely fails (0.0% across all benchmarks)—Hybrid Actor achieves remarkable accuracy across the board: 70.0% on GSM8K, 72.7% on MetaMath, 47.2% on PRM12K, and 6.9% on LIMO itself. These results substantially surpass both SFT+RL and SFT, which often collapse under sparse reward: e.g., SFT+RL yields only 28.7% on GSM8K and 48.7% on MetaMath, while SFT achieves 16.3% and 17.3% respectively. Similarly, when trained on HiTab, Hybrid attains the highest in-domain accuracy of 64.2%, outperforming both SFT+RL (60.7%) and SFT (5.5%) by large margins. Even on out-of-domain tasks like MetaMath and PRM12K, Hybrid maintains competitive performance, demonstrating robust transferability. These results underscore the advantages of Hybrid Actor in leveraging supervised demonstrations while maintaining the exploratory flexibility of reinforcement learning. By integrating both signals throughout training, it avoids brittle mode transitions and achieves consistently strong performance across domains—particularly in sparse-reward environments where traditional pipelines often fail to generalize.
4.4 Generalization Study
Table 3: Exact-match accuracy (%) on GSM8K, PRM12K, and LIMO when trained on LIMO using Qwen2.5-1.5B under PPO and GRPO. Vanilla baselines are shown in gray; gains/losses are highlighted in green/red. Hybrid refers to the Hybrid Actor.
Optimizer Method GSM8K PRM12K LIMO
PPO PPO 52.5 29.9 4.9
SFT 59.7 26.7 1.2
SFT+PPO 70.3 44.4 8.5
PPO Hybrid 67.7 48.6 11.6
GRPO GRPO 0.0 0.0 0.0
SFT 59.7 26.7 1.2
SFT+GRPO 28.7 19.3 3.0
GRPO Hybrid 70.0 47.2 6.9
The Hybrid Actor enables stable, transferable learning across different RL algorithms To assess generalization across RL algorithms in sparse-reward settings, we evaluate our Hybrid Actor under both PPO and GRPO using the same model architecture (Qwen2.5-1.5B) and training data (LIMO). As shown in Table 3, Hybrid consistently outperforms both vanilla RL and SFT+RL baselines. Under PPO, Hybrid Actor achieves +18.7% and +6.7% improvements over vanilla PPO on PRM12K and LIMO, respectively. Compared to the stronger SFT+PPO pipeline, Hybrid still yields notable gains: +4.2% on PRM12K and +3.1% on LIMO. Under GRPO, Hybrid delivers even larger margins: +70.0% on GSM8K, +47.2% on PRM12K, and +6.9% on LIMO compared to the vanilla GRPO baseline, which fails to learn effectively. Even when compared to SFT+GRPO, Hybrid achieves consistent gains of +41.3%, +27.9%, and +3.9% on GSM8K, PRM12K, and LIMO, respectively.
Table 4: Exact-match accuracy (%) on GSM8K, PRM12K, HiTab, and LIMO when trained on HiTab using PPO and Hybrid across model scales. PPO serves as baseline (gray); improvements/degradations from Hybrid are shown in green/red. Hybrid refers to the Hybrid Actor.
Model Method GSM8K PRM12K HiTab LIMO
Qwen2.5 1.5B PPO 46.9 30.4 0.1 5.5
Hybrid 43.1 30.6 0.5 5.5
Qwen2.5 3B PPO 68.5 49.1 16.3 12.8
Hybrid 72.1 55.3 53.7 15.2
Qwen2.5 7B PPO 86.6 59.6 65.5 21.3
Hybrid 83.8 63.7 76.4 28.1
LLaMA3.2 1B PPO 41.6 31.9 0.8 6.7
Hybrid 40.5 25.1 0.1 7.3
LLaMA3.2 3B PPO 74.2 53.2 48.4 20.1
Hybrid 74.2 53.8 61.9 23.2
LLaMA3.1 8B PPO 82.3 55.3 68.8 20.1
Hybrid 82.0 54.6 77.7 20.1
The Hybrid Actor generalizes across model scales and architectures.Table 4 presents exact-match accuracy across four datasets when models are trained on HiTab. For smaller models like Qwen2.5-1.5B and LLaMA3.2-1B, Hybrid yields mixed results, with minor regressions on GSM8K and PRM12K. However, starting from the 3B scale, Hybrid begins to outperform PPO. For example, on Qwen2.5-3B, Hybrid achieves +6.2% on PRM12K and a substantial +37.4% on HiTab. The gains persist on Qwen2.5-7B, with Hybrid reaching +10.9% improvement on HiTab and +6.8% on LIMO. Similarly, for LLaMA3.2-3B, Hybrid improves HiTab accuracy from 48.4% to 61.9%. These results indicate that the advantages of Hybrid training persist across both model scales and architectures (e.g., Qwen and LLaMA), demonstrating its robustness and effectiveness in sparse-reward environments.
4.5 Empirical Study on K-step Training Parameters for Adaptive Switch
In this section, we discuss the parameter settings of Adaptive Switch in SuperRL. In the Appendix C.8, we study the impact of different initialization values of $\sigma_{pg}$ and $\sigma_{sft}$ for Hybrid Actor. Here, we focus on the appropriate values for the parameters related to the K-step training.
Figure 2: Average Reward Values on Different Benchmarks (0-10, 20-30, 40-50 Steps)
Figure 3: Count of Reward Increases on Different Benchmarks (0-10, 0-30, 0-50 Steps)
We propose a reward-based switching strategy to assess the effectiveness of RL during early training for Adaptive Switch. If early rewards are sparse—characterized by stagnant or low rewards—the system transitions to a hybrid training mode, improving learning stability in sparse-reward settings.
[…]
Table 7 in the Appendix illustrates early reward dynamics during the first 10 steps of GRPO training with Qwen2.5-1.5B on GSM8K and HiTab. On GSM8K, the initially low average reward ($avg\_reward$) increases rapidly, reaching 0.4894 by step 10. This indicates that standard RL training progresses effectively, and no switching is required. In contrast, on HiTab, both the number of reward increases ($increase\_num$) and the recent average reward ($recent\_avg\_reward$) remain near zero, signaling training stagnation and prompting a mode switch.
Figures 3 and 3 provide further insight by plotting the average reward and cumulative count of reward increases across early training steps for Qwen2.5-1.5B on multiple benchmarks. As shown, easier datasets such as GSM8K and MetaMath exhibit steady increases in both metrics, while harder datasets like LIMO and HiTab show little to no improvement. Based on these observations, we empirically set the switch parameters as follows: for batch sizes larger than 32, we use $k=10$, $m=10$, $increase\_threshold=3$, and $recent\_avg\_reward=0.1$; otherwise, we set $k=50$, $m=10$, $increase\_threshold=20$, and $recent\_avg\_reward=0.2$. Full details of the experimental setup are provided in Appendix C.8.

### v2

@@ 4.1 Experimental Setup
We evaluate SuperRL across a diverse suite of reasoning benchmarks that vary in structure, difficulty, and reward sparsity. Our evaluation spans both dense-reward tasks (e.g., GSM8K, MetaMath) and sparse-reward tasks (e.g., OpenR1, PRM12K). In all settings, SuperRL consistently outperforms strong baselines, including pure RL and the conventional two-stage SFT+RL pipeline.
In dense regimes, SuperRL performs comparably or better than RL, suggesting that the SFT fallback is rarely triggered and thus incurs no degradation. In sparse regimes, where rollouts frequently fail, SuperRL activates its supervised fallback, leading to significant gains in success rate, generalization, and training stability.
To gain deeper insights into training dynamics, we track the proportion of samples that trigger RL versus SFT over time. On sparse tasks, SFT is used frequently at the beginning to bootstrap learning and is phased out as the policy improves. In contrast, dense tasks rarely invoke SFT. This per-instance adaptivity balances exploration and supervision effectively, reducing KL divergence variance and promoting stable optimization.
We further ablate several design variants: (1) Advantage-based gating, which replaces reward-checking with advantage-checking; (2) Hybrid-Log-Sigma, which softly blends RL and SFT losses using learned uncertainty weights; and (3) Fixed-schedule hybrids with manually tuned schedules. Although these variants yield some improvements over naïve RL, they require extra tuning or introduce rigidity. In contrast, SuperRL offers a simple, hyperparameter-free, and plug-and-play switching mechanism that delivers stronger and more stable performance.
Table 1: Cross-dataset generalization accuracy (%) of different training methods. Each row block corresponds to a model trained on a specific dataset (e.g., GSM8K, Metamath, etc.), while columns report the zero-shot accuracy on various held-out test sets. The baseline (Vanilla RL) performance is shown in the first row of each block and shaded in gray. Colored cells indicate the relative change compared to this baseline: green for improvement and red for degradation, with darker shades representing larger absolute differences. This visualization highlights the generalization patterns of each training method across diverse reasoning tasks.
Training Method GSM8K Metamath PRM12K LIMO OpenR1 AIME24 AIME25
GSM8K RL 72.29 72.05 33.18 23.58 14.96 2.92 1.10
SFT 58.00 64.29 22.86 0.00 7.14 0.00 0.00
SFT+RL 71.45 74.56 32.14 18.38 14.58 1.45 0.44
SuperRL 78.86 77.02 36.96 53.46 16.11 4.10 1.81
Metamath RL 76.21 81.03 38.16 41.63 15.01 6.67 2.05
SFT 69.43 74.29 22.14 14.29 12.86 0.00 3.33
SFT+RL 71.44 72.14 32.14 9.74 12.31 2.99 0.64
SuperRL 77.00 84.29 39.09 48.08 15.73 4.73 1.81
LIMO RL 73.28 76.66 37.26 44.13 16.82 3.79 1.66
SFT 39.14 35.00 21.43 0.00 10.71 0.00 3.33
SFT+RL 77.44 76.72 35.75 34.88 15.24 2.85 6.67
SuperRL 73.28 66.43 35.64 36.99 14.92 3.99 1.14
PRM12K RL 64.61 74.40 44.48 25.14 19.65 2.59 2.09
SFT 61.43 57.86 22.86 0.00 11.43 0.00 0.00
SFT+RL 67.56 80.46 42.99 28.36 16.61 2.71 1.21
SuperRL 70.95 81.45 48.10 39.57 19.35 4.28 2.51
OpenR1 RL 56.98 67.78 42.10 37.42 17.40 2.91 0.99
SFT 65.00 40.71 28.57 28.57 11.43 3.33 0.00
SFT+RL 66.77 67.53 30.66 26.98 14.93 3.04 1.42
SuperRL 73.68 77.98 37.21 33.62 19.68 3.55 1.84
4.2 Cross-Dataset Generalization Analysis
Table 2: Cross-dataset test accuracy (%) of models trained on Hitab dataset. Colors indicate change relative to RL (top row): green = higher accuracy, red = lower accuracy; deeper shades reflect larger deltas.
Training Method GSM8K Metamath PRM12K LIMO OpenR1 AIME24 AIME25 Hitab
Hitab RL 72.05 44.21 26.63 10.94 15.26 2.99 0.57 30.03
SFT 64.29 55.00 33.57 28.57 11.43 0.00 0.00 33.33
SFT+RL 74.56 72.53 35.91 36.93 16.65 4.41 4.11 38.64
SuperRL 77.40 34.29 29.29 19.43 12.45 3.33 2.08 18.47
Key Observations
Across the $5{\times}7=35$ pairs in Table 1, SuperRL improves upon RL in 24 cases, is worse in 10, and ties once. [^1] A cell is counted as an improvement when the SuperRL entry strictly exceeds the corresponding RL entry. Gains are largest when the source dataset itself supplies either dense rewards (e.g., GSM8K, MetaMath) or moderately sparse but still informative feedback (PRM12K, OpenR1). In these settings, reward–gated switching allows exploration-driven RL updates while still injecting high-quality supervision on difficult examples, yielding improvements of up to +29.9 pp.
Insights from LIMO
When trained on LIMO, SuperRL underperforms RL—a divergence attributable to both the dataset’s limited size and its deliberate construction. Specifically, LIMO contains only 817 examples, all manually filtered to emphasize high reasoning complexity. This results in sparse rewards and low pattern diversity: successful trajectories are few and narrowly distributed. In such settings, RL can overfit to these recurring patterns, achieving higher in-domain accuracy. However, SuperRL’s fallback to SFT introduces broader—but diffuse—supervision that struggles to provide useful learning signals when data is both scarce and hard. Consequently, SuperRL cannot effectively learn from LIMO alone. Nevertheless, SuperRL demonstrates strong generalization despite its poor in-domain performance. When trained on more diverse and densely supervised datasets such as GSM8K, it achieves surprisingly strong results on LIMO and other hard transfer benchmarks. This contrast highlights the critical role of training data: it not only affirms the limitations of LIMO’s narrow construction but also reinforces SuperRL’s ability to generalize when given sufficient and well-distributed supervision. The findings suggest that the model’s underperformance is more reflective of dataset-specific bottlenecks than of architectural limitations.
Results in Table 2 show a clear contrast with earlier trends: the two-stage SFT+RL achieves the best performance on Hitab, while SuperRL falls behind even vanilla RL. Hitab is a challenging, cross-domain benchmark focused on table reasoning, where rewards are binary and high-quality demonstrations—collected via DeepSeek R1 distillation and RFT-style sampling—are abundant. In this setting, a dedicated SFT phase proves critical: it allows the model to absorb diverse, domain-specific reasoning patterns before engaging in exploration. SuperRL, by contrast, switches prematurely to RL as soon as a single successful trajectory is found, which may truncate the supervision phase too early and miss the chance to learn stable table reasoning heuristics. This issue is especially pronounced in cross-domain settings like Hitab, where general reasoning strategies learned via minimal supervision may not transfer well. Nevertheless, SuperRL still exhibits strong generalization ability across datasets. Despite its underperformance on Hitab, it maintains competitive transfer performance on other challenging benchmarks—highlighting that its fallback supervision discourages brittle memorization and fosters more broadly applicable reasoning strategies.
Practical take-aways.
  • Reward-gated switching is generally effective: it provides a plug-and-play improvement over RL with no extra hyperparameters in most transfer scenarios.
  • Dataset properties matter: highly filtered, ultra-sparse sources such as LIMO may require additional stabilisation; heterogeneous, demonstration-rich benchmarks such as Hitab benefit from longer SFT exposure.
  • Future work: learn a data-driven gating schedule that adapts to both reward sparsity and demonstration density, thereby unifying the advantages of SuperRL and two-stage SFT+RL.
4.3 Adaptive Behaviors During Training
Beyond accuracy and generalization, we also assess SuperRL in terms of training stability and computational cost. These dimensions are critical for real-world deployment, especially in settings where reward sparsity and rollout variance can induce unstable gradients and excessive tuning overhead.
Entropy Stabilization Patterns.
Table 3: Entropy statistics of RL and SuperRL across datasets. Green cells indicate improvements over RL (i.e., lower entropy), with darker shades representing greater reductions.
Method & Dataset Max Min Range Variance
RL (LIMO) 5.794 0.006 5.788 0.831
SuperRL (LIMO) 4.698 0.009 4.689 0.490
RL (GSM8K) 4.850 0.004 4.846 0.171
SuperRL (GSM8K) 4.875 0.006 4.869 0.134
RL (Metamath) 4.863 0.016 4.847 0.424
SuperRL (Metamath) 5.417 0.020 5.397 0.405
RL (OpenR1) 4.978 0.044 4.934 0.708
SuperRL (OpenR1) 5.208 0.034 5.174 0.695
RL (HITAB) 6.002 0.000 6.002 0.363
SuperRL (HITAB) 5.376 0.006 5.370 0.186
To evaluate training stability, we analyze the entropy statistics of RL and SuperRL across multiple datasets, as shown in Table 3. Entropy quantifies the uncertainty of the policy during exploration; high variance or wide ranges often indicate unstable updates, while excessively low entropy can suppress exploration. A well-calibrated entropy profile reflects both stable learning and sufficient exploration capacity.
Across most datasets, SuperRL achieves a meaningful reduction in entropy variance and range compared to vanilla RL, suggesting more stable and controlled optimization. For instance, in GSM8K and Metamath, SuperRL consistently narrows the entropy range and lowers the variance, leading to smoother convergence without sacrificing exploration. These improvements reflect the benefit of selectively falling back to supervised fine-tuning in failure cases.
However, on LIMO and HITAB, we observe an over-suppression of entropy dynamics: while SuperRL lowers the entropy variance (e.g., from $0.831$ to $0.490$ on LIMO), it also curtails the range and maximum entropy to a degree that may hinder sufficient exploration. This suggests that frequent fallback to SFT may anchor the policy too tightly to demonstrative behaviors, limiting the diversity of trajectories explored. Consequently, the model may prematurely exploit suboptimal behaviors, leading to diminished performance on exploration-heavy tasks.
These findings highlight that entropy stabilization is a double-edged sword: while reducing variance helps avoid erratic learning, overly dampened entropy signals can impair policy expressiveness. SuperRL performs best when the fallback mechanism maintains a dynamic balance between policy stability and exploratory diversity. In aggregate, its adaptive optimization strategy yields more reliable convergence across diverse reward landscapes, though careful tuning may be necessary to avoid under-exploration in sparse regimes.
SFT Trigger Patterns.
Figure 2: Analysis of SFT Activation Patterns and Efficiency. Left: The histogram shows the count of supervised fine-tuning (SFT) activations per 100 training steps for two datasets: Metamath and LIMO. Most SFT activations in both cases occur in the early phase of training (0–100 steps), with LIMO exhibiting a denser activation pattern overall. Right: A comparison of SFT activation efficiency metrics across datasets. LIMO shows longer average activation lengths, shorter intervals between activations, and a higher activation density (segments with SFT triggered per 500 steps), indicating more frequent and sustained reliance on SFT throughout training.
We observe notable differences in how frequently and efficiently supervised fine-tuning (SFT) is triggered across datasets. As shown in Figure 2, LIMO triggers SFT much more frequently than Metamath, especially in the early stages of training (0–100 steps), suggesting that it encounters a higher proportion of hard instances requiring fallback supervision. Additionally, the right panel highlights that LIMO exhibits longer average SFT activation lengths and denser activation patterns, with shorter intervals between activations. These trends imply that the model struggles to sustain reward-guided learning for longer stretches and must rely more consistently on offline signals. In contrast, Metamath shows more sparsely distributed SFT usage, indicating more stable or successful online exploration with less dependence on supervised corrections once initial learning has occurred.
Importantly, across both datasets, we observe a consistent trend: the reliance on SFT signals decreases as training progresses. This pattern highlights the adaptive nature of our method—SFT is heavily used in early stages to guide learning when the model lacks competency, but gradually fades out as the policy improves and the model learns to solve problems correctly on its own. This transition reflects a shift from supervision-driven learning to reward-driven generalization, demonstrating the effectiveness of our hybrid training framework.
Table 4 provides a quantitative summary of average scores and SFT fallback usage across datasets. A clear negative correlation emerges between the SuperRL average score and total SFT trigger count (r=0.86r=-0.86, p=0.028< 0.05p=0.028\ <\ 0.05), suggesting that fallback is most critical on tasks where standard RL struggles to achieve high scores. This trend is even more pronounced when comparing score gains: in datasets with high SFT fallback (e.g., PRM12K, OpenR1), SuperRL significantly outperforms RL. Conversely, in easier datasets like GSM8K or Metamath, where scores are relatively high and SFT is rarely needed, the margin is small or even reversed.
This analysis underscores the adaptive value of SuperRL’s fallback mechanism. When RL alone fails to make progress—either due to sparse rewards or brittle dynamics—SFT provides essential guidance. Rather than rely on static supervision, SuperRL selectively triggers imitation based on rollout outcomes, offering a robust and flexible solution for navigating varying task difficulty. The heterogeneous SFT usage patterns thus reflect not only task complexity but also the system’s capacity to maintain learning momentum across regimes.
Interestingly, when analyzing the relationship between RL average score and SFT trigger count, we observe a similar but weaker negative correlation (r=0.64r=-0.64, $p\approx 0.171$). While the trend aligns with the intuition that lower RL performance invites more SFT fallback, the lack of statistical significance suggests that raw RL scores alone are not as strong a predictor of fallback reliance. This is likely because SuperRL selectively invokes SFT based on rollout success rather than static task characteristics—meaning that even if RL achieves moderately good average scores on some tasks, it may still benefit from occasional fallback in rare failure cases.
In contrast, the SuperRL score shows a much stronger and statistically significant correlation with SFT usage, reinforcing the idea that SuperRL dynamically calibrates its reliance on supervision to support progress where RL plateaus. Taken together, these results highlight that SFT fallback not only enhances performance but also reflects the model’s internal learning difficulty, offering a principled mechanism to bridge the gap between exploration and convergence.
Table 4: Average score and SFT-triggered update count during training. Higher SFT usage correlates with harder tasks and greater gains in average score.
Dataset RL Score SuperRL Score SFT Trigger Count
OpenR1 0.526 0.619 728
PRM12K 0.268 0.590 484
HITAB 0.645 0.665 270
LIMO 0.500 0.688 247
Metamath 0.715 0.700 40
GSM8K 0.687 0.758 19
4.4 Hybrid Variants and Insights
We conclude by comparing SuperRL with two hybrid variants — Hybrid-Advantage-Gated and Hybrid-Log-Sigma, each relaxing the original reward-gated switch in different ways. Table 5 presents the cross-dataset generalization performance under consistent training datasets.
Training Set Method GSM8K Metamath PRM12K LIMO OpenR1 AIME2024 AIME2025
GSM8K SuperRL 78.86 77.02 36.96 53.46 16.11 4.10 1.81
Adv-Gated 76.43 77.93 37.48 34.73 15.50 3.46 2.20
Log-Sigma 73.88 52.72 37.60 24.18 14.29 2.95 1.22
Metamath SuperRL 77.00 84.29 39.09 48.08 15.73 4.73 1.81
Adv-Gated 74.77 84.29 38.26 43.27 15.86 4.21 2.87
Log-Sigma 71.95 77.71 32.60 14.80 14.21 2.50 1.02
LIMO SuperRL 73.28 66.43 35.64 36.99 14.92 3.99 1.14
Adv-Gated 71.20 77.39 36.98 37.54 15.24 4.63 6.67
Log-Sigma 70.55 71.86 34.21 28.96 14.66 3.63 2.36
PRM12K SuperRL 70.95 81.45 48.10 39.57 19.35 4.28 2.51
Adv-Gated 74.90 81.18 51.00 40.87 20.17 2.26 2.11
Log-Sigma 60.97 46.98 42.94 18.06 15.87 1.51 1.86
OpenR1 SuperRL 73.68 77.98 37.21 33.62 19.68 3.55 1.84
Adv-Gated 14.28 29.61 28.57 14.11 10.23 0.56 0.74
Log-Sigma 10.40 27.18 27.53 14.29 12.18 0.18 1.42
Hitab SuperRL 77.40 34.29 29.29 19.43 12.45 3.33 2.08
Adv-Gated 78.09 38.62 29.29 9.76 13.51 5.54 0.64
Log-Sigma 61.08 27.48 29.02 15.78 12.18 3.05 1.09
Table 5: Cross-dataset test accuracy (%) of SuperRL and its variants. Adv-Gated uses advantage-based gating; Log-Sigma softly blends RL and SFT losses with uncertainty-weighted averaging. Color indicates delta relative to SuperRL: green = higher, red = lower.
SuperRL Outperforms in the Majority.
Despite its localized shortcomings on LIMO and Hitab, SuperRL consistently outperforms both Hybrid-Adv-Gated and Hybrid-Log-Sigma in the vast majority of settings. As shown in Table 5, SuperRL achieves higher or equal accuracy in 28 out of 42 train-test pairs compared to the advantage-gated variant, and dominates the uncertainty-based hybrid in nearly every case. This pattern holds especially true when the training dataset offers a reasonable mix of reward sparsity and demonstration richness—conditions under which SuperRL’s reward-gated fallback efficiently balances exploration and imitation.
These results reaffirm the core design philosophy behind SuperRL: by using a simple, deterministic switch based on actual rollout outcomes, it avoids overfitting to noisy shaping signals while also bypassing the need for complex auxiliary mechanisms. In contrast, advantage-based gating introduces extra learning dynamics that may require careful calibration, while soft blending with uncertainty weights often leads to instability or underperformance. The empirical gap across datasets highlights that SuperRL’s binary gating mechanism—though minimal—is highly effective and broadly applicable.
When SuperRL Falls Behind: The Case of LIMO and Hitab.
While SuperRL generally achieves strong generalization across diverse benchmarks, two notable exceptions emerge: LIMO and Hitab. In both cases, the Advantage-Gated variant surpasses SuperRL, suggesting specific regimes where reward-gated switching may be less effective.
In LIMO, the dataset is both small in size (817 samples) and manually curated for high reasoning difficulty, resulting in extremely sparse rewards and low structural diversity. Under such conditions, SuperRL’s reward-based gate may be too permissive—deeming trajectories with minimal reward as successful, thereby deactivating the fallback to supervision prematurely. The Advantage-Gated method, which relies on learned advantage signals, is more selective: it delays the switch until a trajectory meaningfully improves over the baseline. This added caution helps preserve supervision longer during early training, which proves crucial when reward signals are misleading or insufficient for exploration.
In Hitab, the situation is different but equally revealing. Hitab focuses on table-based reasoning, which presents a different set of challenges compared to math word problems or symbolic reasoning. Furthermore, the dataset is constructed via distillation from DeepSeek R1 followed by RFT-style filtering, resulting in trajectories that reflect model-like patterns rather than diverse human logic. In this context, SuperRL’s reward-triggered switching can be overly aggressive: once it detects any success signal—even if superficial—it disables the SFT path, missing the opportunity to deeply absorb structured table reasoning heuristics. On the other hand, the Advantage-Gated approach defers the switch until the learned advantage becomes significant, thus sustaining supervision longer and capturing the domain-specific inductive biases required for table reasoning.
Interpretation and Value of Reward-Gated Switching.
These failure cases highlight an important limitation: although reward-based gating is efficient and principled in most reasoning tasks, it may struggle when (i) reward signals are noisy and overly sparse (e.g., LIMO), or (ii) task structure demands more sustained exposure to supervision (e.g., table-based reasoning in Hitab).
Nevertheless, SuperRL remains highly competitive overall. Its performance on other hard benchmarks—especially when trained on more general-purpose datasets like GSM8K or PRM12K—shows that the reward-gated fallback mechanism is effective in dynamically routing learning signals based on rollout success, without the need for extra models or hyperparameters.
Takeaway.
While Advantage-Gated may outperform in specific scenarios like LIMO and Hitab, SuperRL offers a broadly applicable, plug-and-play strategy that delivers reliable performance across varied reasoning domains. Future work may explore combining reward- and advantage-based signals or developing more domain-aware switching strategies—but the current design already strikes a strong balance between stability, adaptability, and simplicity.